OpenStudy (.whitedragon.):

Forrest Lumber uses the function S(t) = -105t + 630 to determine the salvage value S(t), in dollars, of a table saw t years after its purchase. How long will it take the saw to depreciate completely? 7 years 5 years 6 years -6 years

geerky42 (geerky42):

For it to depreciate completely, S(t) must be equal to 0

geerky42 (geerky42):

So you have: S(t) = -105t + 630 0 = -105t + 630 Can you find t now?

OpenStudy (.whitedragon.):

no idk

geerky42 (geerky42):

To find t, you need to isolate t. So we have 0 = -105t + 630 We need to do something about 630. What to do?

OpenStudy (.whitedragon.):

idk subtract

OpenStudy (.whitedragon.):

or add

OpenStudy (anonymous):

divide.

OpenStudy (anonymous):

s(t)=-105t+630 -105t=s(t)-630 t=(630-s(t)/105

OpenStudy (anonymous):

t=s(t)-630/105

OpenStudy (anonymous):

divide this and u get the answer.